OpenStudy (anonymous):

Write 4*10-2 as a decimal.

OpenStudy (johnweldon1993):

\[\large 4 \times 10^{-2} = 4 \times \frac{1}{10^2} =\ \frac{4}{10^2} = \frac{4}{100} = ?\]

OpenStudy (anonymous):

do i simplify?

OpenStudy (anonymous):

or turn into decimal actually?

OpenStudy (anonymous):

0.04?

OpenStudy (johnweldon1993):

Yeah decimal...and correct

OpenStudy (anonymous):

thx so much :)

OpenStudy (anonymous):

can you help me with more?